How do you find the valency of chlorine, sulphur and magnesium?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Try to recall that chlorine is an element of group 17, sulphur is an element of group 16 and magnesium is an element of group 2 of modern periodic table. Complete step by step solution: We know that, valency of an element is basically its tendency to accept or lose electrons in order to complete its octet and get stable electronic configuration. It is the minimum number of electrons we have to add or remove in order to completely fill the outermost shell of an element. Mathematically we can say that if the outermost shell of an atom contains 4 or less than 4 electrons, then the valency of an element is equal to the number of electrons present in the outermost shell and if it is greater than 4, then the valency of an element is determined by subtracting the total number of electrons present in outermost shell from 8. Calculation of valency: Valency of chlorine: The electronic configuration of chlorine=2, 8, 7. Since, chlorine has 7 (more than 4) electrons in its outermost shell. Therefore, valency of chlorine = 8-number of electrons in outermost shell =8−7 =1 Valency of Sulphur: The electronic configuration of Sulphur=2, 8,6. Since, Sulphur has 6 (more than 4) electrons in its outermost shell. Therefore, valency of chlorine = 8-number of electrons in outermost shell =8−6 =2. Now, Valency of magnesium: The electronic configuration of Magnesium =2, 8, 2. Since, magnesium has 2 (less than 4) electrons in its outermost shell. Therefore, valency of magnesium= Number of electrons in its outermost shell =2 Note: It should be remembered that on going from left to right along a period, valency first increases from 1 to 4 and then decreases from 4 to 1.